# Knowledge in multi-robot systems: an interplay of dynamics, computation and communication

### v1

@@ 4.2 The Exploration Task in the LCM Model
Consider a model $\mathcal{I}$ with an exploration space $X$, from a $\mathcal{LUMI}$ robot system with full visibility and exploration flooding, and $\text{exp}(U)$ an exploration statement, an arbitrary run $\sigma$ from $\mathcal{I}$, and an arbitrary robot $\mathbf{r}\in\Pi$. If $\mathcal{I},\sigma\vDash\Diamond K_{\mathbf{r}}\text{exp}(U)$, then $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(U)$.
Proof 4.14.
Since $\mathcal{I},\sigma\vDash\Diamond K_{\mathbf{r}}\text{exp}(U)$, then there exists a time $[t_{0}]$ such that $\mathcal{I},\sigma,[t_{0}]\vDash K_{\mathbf{r}}\text{exp}(U)$. Since robots have perfect recall, and exploration statements are stable, $\mathcal{I},\sigma,[t]\vDash K_{\mathbf{r}}\text{exp}(U)$ for any $[t]\geqslant[t_{0}]$.
Let $\mathbf{r}^{\prime}\in\Pi$ be an arbitrary robot. Since we don’t consider crash failures, there is a time $[t]_{\mathbf{r}^{\prime}}\geqslant[t_{0}]$ when $\mathbf{r}^{\prime}$ is able to read the lights of robot $\mathbf{r}$. Let $U^{\prime}$ be the maximal region that $\mathbf{r}$ knows to be explored at time $[t]_{\mathbf{r}^{\prime}}$. Since the robot system is exploration flooding, then robot $\mathbf{r}$ is communicating $\text{exp}(U^{\prime})$ through its lights. Therefore, robot $\mathbf{r}^{\prime}$ learns $\text{exp}(U^{\prime})$, i.e. $K_{\mathbf{r}^{\prime}}\text{exp}(U^{\prime})$. However, recall that $U\subseteq U^{\prime}$. Thus $\mathcal{I},\sigma,[t]_{\mathbf{r}^{\prime}}\vDash K_{\mathbf{r}^{\prime}}\text{exp}(U)$.
Since $\Pi$ is finite, then there is a $[t_{\textrm{max}}]$ such that $\mathcal{I},\sigma,[t_{\textrm{max}}]K_{\mathbf{r}^{\prime\prime}}\text{exp}(U)$ for any robot $\mathbf{r}^{\prime\prime}\in\Pi$. Finally, $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(U)$.
@@
Consider an interpreted system $\mathcal{I}$ with a compact exploration space $X$ such that $\mathcal{I}$ satisfies the exploration liveness condition, and such that the robot system is exploration flooding. Then $\mathcal{I}$ satisfies the cooperative termination property from Definition 4.6, namely, $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$
Proof 4.16.
Since $X$ is compact and satisfies exploration liveness, then by Theorem 4.8, then $\mathcal{I}\vDash\Diamond\text{exp}(X)$. By exploration agency, then $\mathcal{I}\vDash\Diamond D_{\Pi}\text{exp}(X)$. Let $\sigma$ be a fixed run of $\mathcal{I}$, then there exists a time $[t_{0}]$ such that $\mathcal{I},(\sigma,[t_{0}])\vDash D_{\Pi}\text{exp}(X)$. Consider for each robot $\mathbf{r}\in\Pi$, a region $U_{\mathbf{r}}:=\bigcup\{U\subseteq X\;|\;\mathcal{I},\sigma,[t_{0}]\vDash K_{\mathbf{r}}\text{exp}(U)\}$. Note that $K_{\mathbf{r}}\text{exp}(U_{\mathbf{r}})$, and $U_{\mathbf{r}}$ is the maximal region that robot $\mathbf{r}$ knows to be explored. From exploration independenceDefinition 4.5), it follows that $\bigcup_{i=1}^{k}U_{i}=X$.
From Lemma 4.13, it holds that $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}U_{\mathbf{r}}$. Therefore, $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\bigcup_{i=1}^{n}U_{i}$. This shows that $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(X)$.
Let $[t_{1}]$ be such that $\mathcal{I},\sigma,[t_{1}]\vDash E_{\Pi}\text{exp}(X)$. In particular consider an arbitrary robot $\mathbf{r}^{\prime}\in\Pi$. Since $\mathcal{I},\sigma,[t_{1}]\vDash K_{\mathbf{r}}\text{exp}(X)$, then by definition of knowledge, in any other run $\sigma^{\prime}$ and any other time $[t^{\prime}]$ that is indistinguishable from $\sigma$ at time $[t_{1}]$ to $\mathbf{r}^{\prime}$, $\text{exp}(X)$ holds. From Lemma 4.13, it follows that $\mathcal{I},\sigma^{\prime},[t^{\prime}]\vDash\Diamond E_{\Pi}\text{exp}(X)$. Thus, from the definition of the knowledge modality, $\mathcal{I},\sigma,[t_{1}]\vDash K_{\mathbf{r}^{\prime}}\Diamond E_{\Pi}\text{exp}(X)$. Since $\mathbf{r}^{\prime}$ is arbitrary, then $\mathcal{I},\sigma,[t_{1}]\vDash E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$. Therefore $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$. Therefore, $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$.
@@
Proof 4.18.
Note that Cooperative Termination $\Rightarrow$ Selfish termination $\Rightarrow$ Parallel Termination. Finally note that Theorem 4.15 shows that Parallel Termination $\Rightarrow$ Cooperative Termination, which completes the equivalence cycle.
We can extend our result to robot systems with arbitrary communication, nevertheless, we still need to define a liveness condition for the communication regarding exploration.
@@
We say that a model $\mathcal{I}$ with an exploration space $X$ satisfies the communication liveness condition if for any run $\sigma$ of $\mathcal{I}$, any time $[t]$, any robot $\mathbf{r}$, and any exploration statement $p_{u}$, $\mathcal{I},(\sigma,[t])\vDash K_{\mathbf{r}}\text{exp}(U)\Rightarrow\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{exp}(U)$.
@@
Consider a model $\mathcal{I}$ with a compact exploration space $X$ such that $\mathcal{I}$ satisfies the exploration liveness and communication liveness conditions, and such that the robot system is exploration flooding. Then $\mathcal{I}$ satisfies the cooperative termination property from Definition 4.6, namely, $\mathcal{I}\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{exp}(X)$
Proof 4.21.
Note that we can reuse the proof of Theorem 4.15, but using Definition 4.19 instead of Lemma 4.13.
4.3 The Surveillance Task
@@
We will consider a normalized time frame for detection, represented by the closed interval $[0,1]\subset\mathbb{R}$. As with exploration, we consider a compact space $X$, which we will call our surveillance space. Note that the product space of our surveillance space with the normalized time frame, $X\times[0,1]$ is a compact topological space with the product topology, which we will call the surveillance cylinder.
Note that each exploration region observed by the robots in the surveillance also induces an exploration region in the surveillance cylinder by simply taking the time of the observations into account. For instance, if a robot explores a region $U$ of the surveillance space at a time $t$, then it already has explored $U\times t$ within the surveillance cylinder. Furthermore, if the intruder’s speed is known to be bounded by a factor of $f$, then the exploration of $U$ at time $t_{1}$, implies that that at the next time $t_{2}$ at which the robot is active, $(U\setminus B_{f\cdot t}(\partial U)),t)_{t\in[t_{1},t_{2}]}$ is also explored, where $B_{f\cdot t}(\partial U)$ is the ball of radius $f\cdot t$ around the boundary of $U$. This comes from the fact that if the intruder was not observed inside $U$ at time $t_{1}$, then the farthest that the intruder can reach at a time $t\in[t_{1},t_{2}]$ is given by the ball of radius $f\cdot t$ around the boundary of $U$.
For the surveillance task, we will consider a set of propositional atoms consisting of exploration statements of the surveillance cylinder $X\times[0,1]$, namely $\text{P}_{X\times[0,1]}$. Each of these exploration statements represent that a part of the surveillance cylinder was explored and no intruder was found. In addition to $\text{P}_{X\times[0,1]}$, we also consider two additional propositional atoms, FOUND, and SECURE. Thus, we can define a set of propositional atoms for surveillance, denoted by $\text{P}_{Surv(X)}:=\text{P}_{X\times[0,1]}\cup\{\text{FOUND},\text{SECURE}\}$. Furthermore, we consider an epistemic temporal language for surveillance in the same way as we did for exploration, but with $\text{P}_{Surv(X)}$ instead of $\text{P}_{X}$.
@@
The semantics of $\mathcal{L}_{surv}$ is analogous to the one of $\mathcal{L}_{exp}$ in Definition 4.3. Note that the surveillance space is considered to be secure iff there does not exist a path $\xi:[0,1]\rightarrow X$ such that the trace of $\xi$. $tr(\xi):=\{(\xi(t),t)\;|\;t\in[0,1]\}$ is not explored. A naive approach to secure the surveillance cylinder would be to consider the exploration on $X\times[0,1]$. However this is largely impractical, and often unfeasible, as it is equivalent to observing the space $X$ at all times during the time interval $[0,1]$. Instead, it is sufficient to explore a "cut" or section of the surveillance cylinder. For instance, if the surveillance space $X$ was explored completely at a fixed time time, i.e., if the set $X\times\{t\}$ from the surveillance cylinder is explored, then the path of any intruder would intersect $X\times\{t\}$. More generally, if there is a manifold in the surveillance cylinder that has been observed, and can be expressed as the trace of a continuous function $f:X\rightarrow[0,1]$; $\text{tr}(f):=\{(x,f(x))\;|\;x\in X\}$.
@@
Intuitively, this lemma shows that continuous manifolds on $X$ work as a "catching net" that is able to capture the path of any invader. Therefore, in addition to the conditions required for simple exploration ( Definition 4.5), we also consider the following conditions:
@@
Observation Correctness Consider an arbitrary run $\sigma$, and an intruder path $\alpha:[0,1]\rightarrow X$. Let $U$ be a region of the surveillance cylinder $X\times[0,1]$ such that $U\cap\text{tr}(\alpha)\neq\varnothing$. Then $\neg\text{exp}(U)$.
Securement Let $f:X\rightarrow[0,1]$ be any continuous function. Let $U$ be a region of the surveillance cylinder $X\times[0,1]$, if $\text{tr}(f)\subseteq U$, then $\text{exp}(U)\rightarrow\text{SECURE}$.
Completion condition $\mathcal{I}\vDash(\text{SECURE}\implies\Box\neg\text{FOUND})\wedge(\text{FOUND}\implies\Box\neg\text{SECURE})$.
Detection agency means that the FOUND atom only holds if an intruder has already been detected by a robot. Observation correctness means that a region in the surveillance cylinder is not secure if an intruder is moving through a space-time point in the region, i.e., exploration regions in this context refer only to regions where no intruder has been found. Securement captures the idea from Lemma 4.23 that shows that it is sufficient to explore the trace of a continuous manifold in $X$. Completion condition states that if part of the surveillance cylinder was explored then the intruder was not found (and never will), and if the intruder was found then the surveillance cylinder was not secure (and never will be). Also note that if the trace of the manifold has even a single discontinuity point, then this may be enough for an intruder to evade detection.
@@
Consider a robot system and a model $\mathcal{I}$ consistent with Definition 4.5 with respect to its surveillance cylinder, that satisfies Definition 4.25. Let $\sigma$ be a run of $\mathcal{I}$. We say that the model solves the surveillance task if $\mathcal{I},\sigma\vDash\Diamond\text{FOUND}\vee\Diamond\text{SECURE}$.
Note that by Completion condition (Definition 4.25) SECURE and FOUND atoms are mutually exclusive in a run.
@@
Parallel Termination $\mathcal{I},\sigma\vDash\Diamond D_{\Pi}\text{FOUND}\vee\Diamond D_{\Pi}\text{SECURE}$
Selfish Termination $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{FOUND}\vee\Diamond E_{\Pi}\text{SECURE}$
Cooperative Termination $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{FOUND}\vee\Diamond E_{\Pi}\Diamond E_{\Pi}\text{SECURE}$
Since Lemma 4.23 shows that exploring a continuous manifold of the surveillance cylinder is enough to detect an intruder, it is also enough to fully secure the space, thus we can replace the exploration liveness condition to a manifold instance.
@@
Consider a robot system that satisfies the simple exploration conditions, Definition 4.5, for the surveillance cylinder $X\times[0,1]$, the surveillance conditions, Definition 4.25, and the manifold surveillance liveness condition, Definition 4.28. Then the surveillance task is solved at $\sigma$.
Proof 4.30.
Follows from Lemma 4.23, and Theorem 4.10.
@@
Proof 4.32.
Let $\sigma$ be a run of $\mathcal{I}$, from Theorem 4.29, $\mathcal{I},\sigma\vDash\Diamond\text{FOUND}\vee\Diamond\text{SECURE}$. Recall that by Completion condition (Definition 4.25) SECURE and FOUND are mutually exclusive in a run. First, let us assume that $\mathcal{I},\sigma\vDash\Diamond\text{FOUND}$. From Definition 4.25, there must exist a robot $\mathbf{r}_{i}$ such that finds an intruder at a time $t_{i}$. Since $\mathbf{r}_{i}$ detects the intruder, it holds that $\mathcal{I},\sigma,t_{i}\vDash K_{\mathbf{r}_{i}}\text{FOUND}$. From Definition 4.19, it holds that $\mathcal{I}\sigma\vDash\Diamond E_{\Pi}\text{FOUND}$. In particular, in all runs $\sigma^{\prime}$ that are indistinguishable to $\mathbf{r}_{i}$, $\mathcal{I},\sigma^{\prime}\vDash\Diamond E_{\Pi}\text{FOUND}$. Therefore, $\mathcal{I},\sigma,t_{i}\vDash K_{\mathbf{r}_{i}}\Diamond E_{\Pi}\text{FOUND}$. Since $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\text{FOUND}$, then we can repeat the previous argument for any other robot. Therefore $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{FOUND}$.
Now, assume that $\mathcal{I},\sigma\vDash\Diamond\text{SECURE}$ holds. Then, it follows that no intruder was found. From Definition 4.28, it holds that there exists a continuous manifold $f_{\sigma}:X\rightarrow[0,1]$, such that $\mathcal{I},\sigma\vDash\Diamond p_{\text{tr}(f_{\sigma})}$. Note that the conditions for Theorem 4.20 hold, therefore, $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}p_{\text{tr}(f_{\sigma})}$. From the securement condition from Definition 4.25, it follows that $\mathcal{I},\sigma\vDash\Diamond E_{\Pi}\Diamond E_{\Pi}\text{SECURE}$.

### v2

@@ 4.2 Application to Exploration
Consider a model $\mathcal{I}$ with an exploration space $X$, from a $\mathcal{LUMI}$ robot system with full visibility and exploration flooding, and $\text{sp}(U)$ an space statement, an arbitrary run $\sigma$ from $\mathcal{I}$, and an arbitrary robot $\mathbf{r}\in\Pi$. If $\mathcal{I},\sigma\models\Diamond K_{\mathbf{r}}\text{sp}(U)$, then $\mathcal{I},\sigma\models\Diamond E_{\Pi}\text{sp}(U)$.
Proof 4.13.
Since $\mathcal{I},\sigma\models\Diamond K_{\mathbf{r}}\text{sp}(U)$, then there exists a time $t_{0}]$ such that $\mathcal{I},\sigma,[t_{0}]\models K_{\mathbf{r}}\text{sp}(U)$. Since robots have perfect recall, and space statements are stable, $\mathcal{I},\sigma,[t]\models K_{\mathbf{r}}\text{sp}(U)$ for any $t]\geqslant[t_{0}]$.
Let $\mathbf{r}^{\prime}\in\Pi$ be an arbitrary robot. Since we don’t consider crash failures, there is a time $t]_{\mathbf{r}^{\prime}}\geqslant[t_{0}]$ when $\mathbf{r}^{\prime}$ is able to read the lights of robot $\mathbf{r}$. Let $U^{\prime}$ be the maximal region that $\mathbf{r}$ knows to be explored at time $t]_{\mathbf{r}^{\prime}}$. Since the robot system is exploration flooding, then robot $\mathbf{r}$ is communicating $\text{sp}(U^{\prime})$ through its lights. Therefore, robot $\mathbf{r}^{\prime}$ learns $\text{sp}(U^{\prime})$, i.e. $K_{\mathbf{r}^{\prime}}\text{sp}(U^{\prime})$. However, recall that $U\subseteq U^{\prime}$. Thus $\mathcal{I},\sigma,[t]_{\mathbf{r}^{\prime}}\models K_{\mathbf{r}^{\prime}}\text{sp}(U)$.
Since $\Pi$ is finite, then there is a $t_{\textrm{max}}]$ such that $\mathcal{I},\sigma,[t_{\textrm{max}}]K_{\mathbf{r}^{\prime\prime}}\text{sp}(U)$ for any robot $\mathbf{r}^{\prime\prime}\in\Pi$. Finally, $\mathcal{I},\sigma\models\Diamond E_{\Pi}\text{sp}(U)$.
@@
Consider an interpreted system $\mathcal{I}$ with a compact exploration space $X$ such that $\mathcal{I}$ satisfies the exploration liveness condition, and such that the robot system is exploration flooding. Then $\mathcal{I}$ satisfies the cooperative termination property from Definition˜4.5, namely, $\mathcal{I}\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{sp}(X)$
Proof 4.15.
Since $X$ is compact and satisfies exploration liveness, then by Theorem˜4.7, then $\mathcal{I}\models\Diamond\text{sp}(X)$. By exploration agency, then $\mathcal{I}\models\Diamond D_{\Pi}\text{sp}(X)$. Let $\sigma$ be a fixed run of $\mathcal{I}$, then there exists a time $t_{0}]$ such that $\mathcal{I},(\sigma,[t_{0}])\models D_{\Pi}\text{sp}(X)$. Consider for each robot $\mathbf{r}\in\Pi$, a region $U_{\mathbf{r}}:=\bigcup\{U\subseteq X\;|\;\mathcal{I},\sigma,[t_{0}]\models K_{\mathbf{r}}\text{sp}(U)\}$. Note that $K_{\mathbf{r}}\text{sp}(U_{\mathbf{r}})$, and $U_{\mathbf{r}}$ is the maximal region that robot $\mathbf{r}$ knows to be explored. From exploration independence( Definition˜4.4), it follows that $\bigcup_{i=1}^{k}U_{i}=X$.
From Lemma˜4.12, it holds that $\mathcal{I},\sigma\models\Diamond E_{\Pi}U_{\mathbf{r}}$. Therefore, $\mathcal{I},\sigma\models\Diamond E_{\Pi}\bigcup_{i=1}^{n}U_{i}$. This shows that $\mathcal{I},\sigma\models\Diamond E_{\Pi}\text{sp}(X)$.
Let $t_{1}]$ be such that $\mathcal{I},\sigma,[t_{1}]\models E_{\Pi}\text{sp}(X)$. In particular consider an arbitrary robot $\mathbf{r}^{\prime}\in\Pi$. Since $\mathcal{I},\sigma,[t_{1}]\models K_{\mathbf{r}}\text{sp}(X)$, then by definition of knowledge, in any other run $\sigma^{\prime}$ and any other time $t^{\prime}]$ that is indistinguishable from $\sigma$ at time $t_{1}]$ to $\mathbf{r}^{\prime}$, $\text{sp}(X)$ holds. From Lemma˜4.12, it follows that $\mathcal{I},\sigma^{\prime},[t^{\prime}]\models\Diamond E_{\Pi}\text{sp}(X)$. Thus, from the definition of the knowledge modality, $\mathcal{I},\sigma,[t_{1}]\models K_{\mathbf{r}^{\prime}}\Diamond E_{\Pi}\text{sp}(X)$. Since $\mathbf{r}^{\prime}$ is arbitrary, then $\mathcal{I},\sigma,[t_{1}]\models E_{\Pi}\Diamond E_{\Pi}\text{sp}(X)$. Therefore $\mathcal{I},\sigma\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{sp}(X)$. Therefore, $\mathcal{I}\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{sp}(X)$.
@@
Proof 4.17.
Note that Cooperative Termination $\Rightarrow$ Selfish termination $\Rightarrow$ Parallel Termination. Finally note that Theorem˜4.14 shows that Parallel Termination $\Rightarrow$ Cooperative Termination, which completes the equivalence cycle.
Our result can be extended to cover arbitrary communication, nevertheless, a communication liveness condition would still be necessary.
@@
We say that a model $\mathcal{I}$ with an exploration space $X$ satisfies the communication liveness condition if for any run $\sigma$ of $\mathcal{I}$, any time $[t]$, any robot $\mathbf{r}$, and any space statement $p_{u}$, $\mathcal{I},(\sigma,[t])\models K_{\mathbf{r}}\text{sp}(U)\Rightarrow\mathcal{I},\sigma\models\Diamond E_{\Pi}\text{sp}(U)$.
@@
Consider a model $\mathcal{I}$ with a compact exploration space $X$ such that $\mathcal{I}$ satisfies the exploration liveness and communication liveness conditions, and such that the robot system is exploration flooding. Then $\mathcal{I}$ satisfies the cooperative termination property from Definition˜4.5, namely, $\mathcal{I}\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{sp}(X)$
Proof 4.20.
Note that we can reuse the proof of Theorem˜4.14, but using Definition˜4.18 instead of Lemma˜4.12.
4.3 Application to Surveillance
@@
We will consider a normalized time frame for detection, represented by the closed interval $0,1]\subset\mathbb{R}$. As with exploration, we consider a compact space $X$, which we will call our surveillance space. Note that the product space of our surveillance space with the normalized time frame, $X\times[0,1]$ is a compact topological space with the product topology, which we will call the surveillance cylinder.
Note that each region observed by the robots in the surveillance also induces a region in the surveillance cylinder by simply taking the time of the observations into account. For instance, if a robot explores a region $U$ of the surveillance space at a time $t$, then it already has explored $U\times t$ within the surveillance cylinder. Furthermore, if the intruder’s speed is known to be bounded by a factor of $f$, then the exploration of $U$ at time $t_{1}$, implies that that at the next time $t_{2}$ at which the robot is active, $(U\setminus B_{f\cdot t}(\partial U)),t)_{t\in[t_{1},t_{2}]}$ is also seen, where $B_{f\cdot t}(\partial U)$ is the ball of radius $f\cdot t$ around the boundary of $U$. This comes from the fact that if the intruder was not observed inside $U$ at time $t_{1}$, then the farthest that the intruder can reach at a time $t\in[t_{1},t_{2}]$ is given by the ball of radius $f\cdot t$ around the boundary of $U$.
For the surveillance task, we will consider a set of propositional atoms consisting of space statements of the surveillance cylinder $X\times[0,1]$, namely $\text{P}_{X\times[0,1]}$. Each of these space statements represent that a part of the surveillance cylinder was explored and no intruder was found. In addition to $\text{P}_{X\times[0,1]}$, we also consider two additional propositional atoms, FOUND, and SECURE. Thus, we can define a set of propositional atoms for surveillance, denoted by $\text{P}_{Surv(X)}:=\text{P}_{X\times[0,1]}\cup\{\text{FOUND},\text{SECURE}\}$. Furthermore, we consider an epistemic temporal language for surveillance in the same way as we did for exploration, but with $\text{P}_{Surv(X)}$ instead of $\text{P}_{X}$.
@@
The semantics of $\mathcal{L}_{surv}$ is analogous to the one of $\mathcal{L}_{space}$ in Definition˜4.3. Note that the surveillance space is considered to be secure iff there does not exist a path $\xi:[0,1]\rightarrow X$ such that the trace of $\xi$. $tr(\xi):=\{(\xi(t),t)\;|\;t\in[0,1]\}$ is not explored. A naive approach to secure the surveillance cylinder would be to consider the exploration on $X\times[0,1]$. However this is largely impractical, and often unfeasible, as it is equivalent to observing the space $X$ at all times during the time interval $[0,1]$. Instead, it is sufficient to explore a “cut” or section of the surveillance cylinder. For instance, if the surveillance space $X$ was explored completely at a fixed time time, i.e., if the set $X\times\{t\}$ from the surveillance cylinder is explored, then the path of any intruder would intersect $X\times\{t\}$. More generally, if there is a manifold in the surveillance cylinder that has been observed, and can be expressed as the trace of a continuous function $f:X\rightarrow[0,1]$; $\text{tr}(f):=\{(x,f(x))\;|\;x\in X\}$.
@@
Intuitively, this lemma shows that continuous manifolds on $X$ work as a “catching net” that is able to capture the path of any invader. Therefore, in addition to the conditions required for simple exploration ( Definition˜4.4), we also consider the following conditions:
@@
Observation Correctness Consider an arbitrary run $\sigma$, and an intruder path $\alpha:[0,1]\rightarrow X$. Let $U$ be a region of the surveillance cylinder $X\times[0,1]$ such that $U\cap\text{tr}(\alpha)\neq\varnothing$. Then $\neg\text{sp}(U)$.
Securement Let $f:X\rightarrow[0,1]$ be any continuous function. Let $U$ be a region of the surveillance cylinder $X\times[0,1]$, if $\text{tr}(f)\subseteq U$, then $\text{sp}(U)\rightarrow\text{SECURE}$.
Completion condition $\mathcal{I}\models(\text{SECURE}\implies\Box\neg\text{FOUND})\wedge(\text{FOUND}\implies\Box\neg\text{SECURE})$.
Detection agency means that the FOUND atom only holds if an intruder has already been detected by a robot. Observation correctness means that a region in the surveillance cylinder is not secure if an intruder is moving through a space-time point in the region, i.e., exploration regions in this context refer only to regions where no intruder has been found. Securement captures the idea from Lemma˜4.22 that shows that it is sufficient to explore the trace of a continuous manifold in $X$. Completion condition states that if part of the surveillance cylinder was explored then the intruder was not found (and never will), and if the intruder was found then the surveillance cylinder was not secure (and never will be). Also note that if the trace of the manifold has even a single discontinuity point, then this may be enough for an intruder to evade detection.
@@
Consider a robot system and a model $\mathcal{I}$ consistent with Definition˜4.4 with respect to its surveillance cylinder, that satisfies Definition˜4.24. Let $\sigma$ be a run of $\mathcal{I}$. We say that the model solves the surveillance task if $\mathcal{I},\sigma\models\Diamond\text{FOUND}\vee\Diamond\text{SECURE}$.
Note that by Completion condition (Definition˜4.24) SECURE and FOUND atoms are mutually exclusive in a run.
@@
Parallel Termination $\mathcal{I},\sigma\models\Diamond D_{\Pi}\text{FOUND}\vee\Diamond D_{\Pi}\text{SECURE}$
Selfish Termination $\mathcal{I},\sigma\models\Diamond E_{\Pi}\text{FOUND}\vee\Diamond E_{\Pi}\text{SECURE}$
Cooperative Termination $\mathcal{I},\sigma\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{FOUND}\vee\Diamond E_{\Pi}\Diamond E_{\Pi}\text{SECURE}$
Since Lemma˜4.22 shows that exploring a continuous manifold of the surveillance cylinder is enough to detect an intruder, it is also enough to fully secure the space, thus we can replace the exploration liveness condition to a manifold instance.
@@
Consider a robot system that satisfies the simple exploration conditions, Definition˜4.4, for the surveillance cylinder $X\times[0,1]$, the surveillance conditions, Definition˜4.24, and the manifold surveillance liveness condition, Definition˜4.27. Then the surveillance task is solved at $\sigma$.
Proof 4.29 (Proof of Theorem˜4.28).
Follows from Lemma˜4.22, and Theorem˜4.9.
@@
Proof 4.31 (Proof of Theorem˜4.30).
Let $\sigma$ be a run of $\mathcal{I}$, from Theorem˜4.28, $\mathcal{I},\sigma\models\Diamond\text{FOUND}\vee\Diamond\text{SECURE}$. Recall that by Completion condition (Definition˜4.24) SECURE and FOUND are mutually exclusive in a run. First, let us assume that $\mathcal{I},\sigma\models\Diamond\text{FOUND}$. From Definition˜4.24, there must exist a robot $\mathbf{r}_{i}$ such that finds an intruder at a time $t_{i}$. Since $\mathbf{r}_{i}$ detects the intruder, it holds that $\mathcal{I},\sigma,t_{i}\models K_{\mathbf{r}_{i}}\text{FOUND}$. From Definition˜4.18, it holds that $\mathcal{I}\sigma\models\Diamond E_{\Pi}\text{FOUND}$. In particular, in all runs $\sigma^{\prime}$ that are indistinguishable to $\mathbf{r}_{i}$, $\mathcal{I},\sigma^{\prime}\models\Diamond E_{\Pi}\text{FOUND}$. Therefore, $\mathcal{I},\sigma,t_{i}\models K_{\mathbf{r}_{i}}\Diamond E_{\Pi}\text{FOUND}$. Since $\mathcal{I},\sigma\models\Diamond E_{\Pi}\text{FOUND}$, then we can repeat the previous argument for any other robot. Therefore $\mathcal{I},\sigma\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{FOUND}$.
Now, assume that $\mathcal{I},\sigma\models\Diamond\text{SECURE}$ holds. Then, it follows that no intruder was found. From Definition˜4.27, it holds that there exists a continuous manifold $f_{\sigma}:X\rightarrow[0,1]$, such that $\mathcal{I},\sigma\models\Diamond p_{\text{tr}(f_{\sigma})}$. Note that the conditions for Theorem˜4.19 hold, therefore, $\mathcal{I},\sigma\models\Diamond E_{\Pi}\Diamond E_{\Pi}p_{\text{tr}(f_{\sigma})}$. From the securement condition from Definition˜4.24, it follows that $\mathcal{I},\sigma\models\Diamond E_{\Pi}\Diamond E_{\Pi}\text{SECURE}$.